# Gromov-Wasserstein Barycenters: The Analysis Problem

### v1

@@ 3.2. GW Barycenters via a Fixed-Point Approach: The Analysis Problem
Algorithm 2 implements the analysis scheme laid out in Theorem 3.6, and we refer the reader to App. C for an analysis of its computational complexity.
@@
13:  $\lambda\leftarrow$ $\arg\min_{\lambda\in\Delta_{S-1}}\lambda^{T}\mathcal{Q}\lambda$
@@ 4. A Gradient Method for Gromov-Wasserstein Barycenters
In [14], given a network $\mathbb{X}$, the authors provide a way of computing directional derivatives of the functional $\mathbb{Y}\longmapsto GW^{2}(\mathbb{X},\mathbb{Y})$ in the direction of vectors in a ‘tangent space’ at $\mathbb{Y}$ [6, 13, 56]. Thus, in Sections 4.1, 4.2, and 4.3 we shift to a different approach that allows to compute gradients of functionals in GW space using a blow-up technique, and thus defining a weaker version of GW barycenters as Karcher means. This culminates in our new contributions to the GW analysis problem (4.4), including Theorem 4.7 and Algorithm 4. Connections between the fixed-point and gradient-based approaches are discussed in Section 4.5.
@@ 4.1. Tangent Space, Exponential Map, Directional Derivatives, and Gradients
As introduced in Section 2, let $\mathcal{GW}$ be the Gromov-Wasserstein space of equivalence classes of networks under weak isomorphism $\sim^{w}$. Thus, two networks $\mathbb{Y}$ and $\mathbb{Y}^{\prime}$ belong to the same class $[\mathbb{Y}]$ if and only if $GW(\mathbb{Y},\mathbb{Y}^{\prime})=0$. In this case, a GW OT plan $\pi$ realizing $GW(\mathbb{Y},\mathbb{Y}^{\prime})=0$ is referred to as a self-coupling. In [14], the authors define a formal tangent structure on the GW space $\mathcal{GW}$ as follows: the tangent space of $\mathcal{GW}$ at a network class $[\mathbb{Y}]$ is given by [^10] If we consider only the space of classes of mm-spaces or gm-spaces under the strong isomorphism defined in Definition 2.3, then each $L^{2}$-space in (25) is restricted to its subspace of symmetric functions.
$\mathrm{Tan}_{[\mathbb{Y}]}=\displaystyle\left(\bigcup_{(Y,\omega_{Y},\mu_{Y})\in[\mathbb{Y}]}L^{2}(Y\times Y,\mu_{Y}\otimes\mu_{Y})\right)/\simeq$ (25)
where the quotient is taken under the following equivalence relation defined by self-couplings: given $v\in L^{2}(Y\times Y,\mu_{Y}\otimes\mu_{Y})$, $v^{\prime}\in L^{2}(Y^{\prime}\times Y^{\prime},\mu_{Y^{\prime}}\otimes\mu_{Y^{\prime}})$, for $(Y,\omega_{Y},\mu_{Y})$, $(Y^{\prime},\omega_{Y^{\prime}},\mu_{Y^{\prime}})$ in $[\mathbb{Y}]$, we say that $v\simeq v^{\prime}$ if and only if there a self-coupling $\pi\in\Pi(\mu_{Y},\mu_{Y^{\prime}})$ such that
@@
We denote the class of $v$ as $[v]$. Each space $\mathrm{Tan}_{[\mathbb{Y}]}$ is equipped with the inner product111111i.e., a well-defined, symmetric, positive semidefinite bilinear form on the quotient space $\mathrm{Tan}_{[\mathbb{Y}]}$, which is modeled on Hilbert spaces.
@@
$\langle A,B\rangle_{L^{2}(Y\times Y,\mu_{Y}\otimes\mu_{Y})}=\sum_{i,j\in[M]}A_{ij}B_{ij}\mathrm{q}_{i}\mathrm{q}_{j}=:\mathrm{tr}_{\mathrm{q}}(A^{T}B),$ (26)
where $\mathrm{tr}_{\mathrm{q}}$ denotes the trace inner product weighted by $\mathrm{q}\otimes\mathrm{q}\in\mathbb{R}^{M\times M}$. [^12] To ensure that this defines a norm, we assume that $\mathrm{q}_{j}>0$ for all $j\in[M]$, meaning that the support of $\mu_{Y}$ consists precisely of the $M$ distinct points $\{y_{j}\}_{j=1}^{M}$, as required in Definition 2.1.
@@
$[v]\longmapsto[(Y,\omega_{Y}+v,\mu_{Y})]\qquad\text{ for }v\in L^{2}(Y\times Y,\mu_{Y}\otimes\mu_{Y}).$ (27)
Using our shorthand notation $(\mathbf{Y},\mathrm{q})\in\mathbb{R}^{M\times M}\times\mathcal{P}_{M}$ to denote a representative of $[\mathbb{Y}]$, (27) can be identified with the map $v\mapsto(\mathbf{Y}+v,\mathrm{q})$ for all $v\in\mathbb{R}^{M\times M}$.
@@
Given a functional over the GW space, $G:\mathcal{GW}\to\mathbb{R}$, its directional derivative in the direction of $[v]\in\mathrm{Tan}_{[\mathbb{Y}]}$ is defined by
$D_{v}G([\mathbb{Y}]):=\lim_{t\searrow 0}\frac{G(Exp_{[\mathbb{Y}]}([tv]))-G([\mathbb{Y}])}{t},$ (28)
and its gradient [^13] If $G:\mathcal{GW}\to\mathbb{R}$ is differentiable, that is, if all the directional derivatives (28) exist, and if, moreover, its gradient at $[\mathbb{Y}]$, as defined via the Riesz representation (29), exists, then it is unique (see [14, Lem. 7] and the references therein). by a vector $\nabla G([\mathbb{Y}])\in\mathrm{Tan}_{[\mathbb{Y}]}$ satisfying
$D_{v}G([\mathbb{Y}])=\langle[v],\nabla G([\mathbb{Y}])\rangle_{\mathrm{Tan}_{[\mathbb{Y}]}},\qquad\forall[v]\in\mathrm{Tan}_{[\mathbb{Y}]}.$ (29)
@@ 4.2. Blow-Ups
In [14], particular representatives of the equivalence classes under weak isomorphisms were considered using a blow-up technique, which we will describe [^14] In a broader sense, this idea is similar to that of the barycentric projection widely used in linear OT [66] in order to get transport maps from OT plans. See also [7] for the GW case.. This blow-up idea avoids the need to fix a size $M$ in advance, as required in the in the approach of [46] and reviewed in this paper when studying the GW synthesis problem (10) in Section 3.1.
@@
Consider finite networks $(X=\{x_{i}\}_{i=1}^{N},\omega_{X},\mu_{X})$ and $(Y=\{y_{j}\}_{j=1}^{M},\omega_{Y},\mu_{Y})$, which we simply denote as $(\mathbf{X},\mathrm{p})\in\mathbb{R}^{N\times N}\times\mathcal{P}_{N}$ and $(\mathbf{Y},\mathrm{q})\in\mathbb{R}^{M\times M}\times\mathcal{P}_{M}$151515(we recall the notation $\omega_{X}(x_{i},x_{i^{\prime}})=\mathbf{X}[i,i^{\prime}]$, $\omega_{Y}(y_{j},y_{j^{\prime}})=\mathbf{Y}[j,j^{\prime}]$, $\mu_{X}=\sum_{i=1}^{N}\mathrm{p}[i]\delta_{x_{i}}$, $\mu_{Y}=\sum_{i=1}^{M}\mathrm{q}[j]\delta_{y_{j}}$). Let $\pi\in\Pi(\mathrm{p},\mathrm{q})$ be a coupling between $\mathrm{p}$ and $\mathrm{q}$. Consider the vectors $u(x)=(u(x_{1}),\dots u(x_{N}))$ and $v(y)=(v(y_{1}),\dots,v(y_{M}))$ given by
@@
    Alignment: One can construct a (possibly enlarged) optimal coupling $\pi_{b}\in\Pi(\mathrm{p}_{b},\mathrm{q}_{b})$ for $GW(\mathbb{X}_{b},\mathbb{Y}_{b})$ that induces an optimal GW-transport map $T:X_{b}\to Y_{b}$, which is a bijection [^16] See [20] for existence of optimal GW maps in the context of mm-spaces under strong isomorphisms., i.e., a permutation of the indices in $[M_{b}]$. As highlighted in [14], the crux of the blow-up construction relies on the observation that while $\mathbb{X}_{b}$ and $\mathbb{Y}_{b}$ maintain weak isomorphisms with $\mathbb{X}$ and $\mathbb{Y}$, respectively, the initial (optimal) transport plan $\pi$ between the original measures $\mu_{X},\mu_{Y}$ supported on the original nodes $X,Y$ (represented the vectors $\mathrm{p},\mathrm{q}$), lifts to a map ${T}$ between the new sets of nodes $X_{b},Y_{b}$ such that $T_{\#}\mu_{X_{b}}=\mu_{Y_{b}}$:
    ${\mathbb{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$      plan      $\scriptstyle{\sim^{w}}$${\mathbb{Y}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\sim^{w}}$${\mathbb{X}_{b}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$      map      ${\mathbb{Y}_{b}}$
    For simplicity, after permuting and relabeling the nodes, if necessary, we will assume that $T$ is the identity map (i.e., we identify $X_{b}=Y_{b}$). Consequently, $\mathrm{p}_{b}=\mathrm{q}_{b}$, and $\pi_{b}$ can be regarded as the diagonal matrix with $\mathrm{q}_{b}$ as its diagonal entries, denoted as $\pi_{b}=\mathrm{diag}(\mathrm{q}_{b})$. See Figure 5 for an illustration. Moreover,
    $GW^{2}(\mathbb{X},\mathbb{Y})$ $=\sum_{j,k\in[M_{b}]}|\mathbf{X}_{b}[j,k]-\mathbf{Y}_{b}[j,k]|^{2}\,\mathrm{q}_{b}[j]\mathrm{q}_{b}[k]=\mathrm{tr}_{\mathrm{q}_{b}}\left((\mathbf{X}_{b}-\mathbf{Y}_{b})^{T}(\mathbf{X}_{b}-\mathbf{Y}_{b})\right).$ (30)
@@
If we consider multiple templates $\{[\mathbb{X}^{s}]\}_{s=1}^{S}$ for $S\geq 1$, given a class $[\mathbb{Y}]$, Algorithm 3 is a general blow-up, which facilitates the alignment of each template representative $\mathbb{X}^{s}=(X^{s},\omega_{X^{s}},\mu_{X^{s}})$ with a representative $\mathbb{Y}=(Y,\omega_{Y},\mu_{Y})$, ensuring that the new spaces $Y_{b}={X^{1}_{b}}=\dots={X^{S}_{b}}$ all contain the same number of nodes ($M_{b}$), share the same mass distribution $\mathrm{q}_{b}$ across the nodes, and the identity map is optimal for each GW problem between $\mathbb{X}^{s}_{b}$ and $\mathbb{Y}_{b}$ ($s\in[S]$). Note that the optimal plans between $\mathbb{X}^{s}_{b},\mathbb{X}^{r}_{b}$, for $s\neq r$, are not necessarily the identity. The network $\mathbb{Y}$ plays a special role in this algorithm, and thus when we apply it to templates $\{[\mathbb{X}^{s}]\}_{s=1}^{S}$, we say that we perform an optimal blow-up with respect to $\mathbb{Y}$.
@@
6:     $\pi$leftarrow optimal plan for $GW((\mathbf{X},\mathrm{p}),(\mathbf{Y},\mathrm{q}))$
@@
Proposition 4.3 (Prop. 8 [14] - Revisited proof in App. B.3).
@@
$G_{\lambda}([\mathbb{Y}]):=\frac{1}{2}\sum_{s=1}^{S}\lambda_{s}GW^{2}(\mathbb{X}^{s},\mathbb{Y}).$ (31)

### v2

@@ 3.2. GW Barycenters via a Fixed-Point Approach: The Analysis Problem
Algorithm 2 implements the analysis scheme laid out in Theorem 3.6, and we refer the reader to Suppl. Mat. C for an analysis of its computational complexity.
Corollary 3.7.
Let $\{(\mathbf{X}^{s},\mathrm{p}^{s})\in\mathbb{R}^{N^{s}\times N^{s}}\times\mathcal{P}_{N^{s}}\}_{s\in[S]}$, $M\in\mathbb{N}$, $(\mathbf{Y},\mathrm{q})\in\mathbb{R}^{M\times M}\times\mathcal{P}_{M}$. In the formulation of the GW analysis problem (13) consider any nonnegative loss function $d_{a}(\cdot,\cdot)$ in $\mathbb{R}^{M\times M}\times\mathbb{R}^{M\times M}$ satisfying the identity of indiscernibles. If $\mathbf{Y}\in\mathrm{Bary}_{M,\mathrm{q}}(\{(\mathbf{X}^{s},\mathrm{p}^{s})\}_{s\in[S]})$, then a weight vector $\lambda_{\mathbf{Y}}\in\Delta_{S-1}$ associated with $\mathbf{Y}$ is given by a solution of the minimization problem (18) provided in Theorem 3.6.
Proof.
If $\mathbf{Y}\in\mathrm{Bary}_{M,\mathrm{q}}(\{(\mathbf{X}^{s},\mathrm{p}^{s})\}_{s\in[S]})$, then it is given by the expression (14) for some weight vector $\in\Delta_{S-1}$. Therefore, for any choice of loss $d_{a}(\cdot,\cdot)$ that is nonnegative and satisfies the identity of indiscernibles, we have $d_{a}(\mathbf{Y},\sum_{s\in[S]}\lambda_{s}\,F(\mathbf{Y},s))=0$ for some $\lambda\in\Delta_{S-1}$ (which is not necessarily unique). If we consider $\lambda_{\mathbf{Y}}$ as a solution of the minimization problem (18) in Theorem 3.6, then, from the proof of such theorem, this is equivalent to have (21), i.e,
(23) $\left\|\mathbf{Y}-\sum_{s\in[S]}(\lambda_{\mathbf{Y}})_{s}\,F(\mathbf{Y},s)\right\|_{\mathrm{Frob}}=\lambda_{\mathbf{Y}}^{T}\mathcal{Q}\lambda_{\mathbf{Y}}=0.$
Since $\|\cdot\|_{\mathrm{Frob}}$ is a norm in the space of $M\times M$ square matrices, (23) implies $\mathbf{Y}=\sum_{s\in[S]}(\lambda_{\mathbf{Y}})_{s}\,F(\mathbf{Y},s)$. So, $d_{a}(\mathbf{Y},\sum_{s\in[S]}(\lambda_{\mathbf{Y}})_{s}\,F(\mathbf{Y},s))=0$, and therefore the non-negativity of $d_{a}$ leads to $\lambda_{\mathbf{Y}}\in\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}d_{a}(\mathbf{Y},\mathbf{Y}_{\lambda})^{2}$ (i.e., the solution $\lambda_{\mathbf{Y}}$ provided in Theorem 3.6 is also a solution for the GW analysis problem (13) with arbitrary divergence $d_{a}$). ∎
Remark 3.8.
Node relabeling corresponds, at the matrix level, to permutation similarity: if $P$ is a permutation matrix and $\mathbf{Y}$ is a dissimilarity matrix, then $P\mathbf{Y}P^{T}$ represents the same object with relabeled nodes. If in (13) we take $d_{a}$ to be the Frobenius norm, then the analysis objective is invariant under such relabeling. Indeed, the Frobenius norm is invariant under permutation similarity (i.e., $\|\mathbf{Y}\|_{\mathrm{Frob}}=\|P\mathbf{Y}P^{T}\|_{\mathrm{Frob}}$ for all permutation matrices $P$). Moreover, the (fixed-size) GW barycenter set is closed under relabeling: if $\mathbf{Y}_{\lambda}$ belongs to the GW barycenter family generated by certain templates, then so does $P\mathbf{Y}_{\lambda}P^{T}$ (which corresponds to the same parameter $\lambda$ as $GW(\mathbf{Y}_{\lambda},\,P\mathbf{Y}_{\lambda}P^{T})=0$). Therefore,
$\min_{\lambda}\|\mathbf{Y}-\mathbf{Y}_{\lambda}\|_{\mathrm{Frob}}=\min_{\lambda}\|P(\mathbf{Y}-\mathbf{Y}_{\lambda})P^{T}\|_{\mathrm{Frob}}=\min_{\lambda}\|P\mathbf{Y}P^{T}-\mathbf{Y}_{\lambda}\|_{\mathrm{Frob}}.$
Moreover, our surrogate analysis problem (16) is also permutation/relabeling invariant for any target $(\mathbf{Y},\mathrm{q})$ (inside or outside the barycenter space) when considering $d_{a}$ as the Frobenius norm. Indeed, let $P$ be an arbitrary permutation that produces $(P\mathbf{Y}P^{T},P\mathrm{q})$ (which encodes the same information as the original shape after relabeling the nodes). If $\pi^{s}\in\Pi(\mathrm{p}^{s},\mathrm{q})$ is an optimal plan for $GW((\mathbf{X}^{s},\mathrm{p}^{s}),(\mathbf{Y},\mathrm{q}))$, it holds that $\pi^{s}P^{T}\in\Pi(\mathrm{p}^{s},P\mathrm{q})$ is optimal for $GW((\mathbf{X}^{s},\mathrm{p}^{s}),(P\mathbf{Y}P^{T},P\mathrm{q}))$. In particular, by using $(PAP^{T})\odot(PBP^{T})=P(A\odot B)P^{T}$ (equivariance of the Hadamard product under simultaneous permutation via conjugation), we have $F(P\mathbf{Y}P^{T},s)=\frac{1}{P\mathrm{q}\mathrm{q}^{T}P^{T}}\odot P(\pi^{s})^{T}\mathbf{X}^{s}\pi^{s}P^{T}=P\left(\frac{1}{\mathrm{q}\mathrm{q}^{T}}\odot(\pi^{s})^{T}\mathbf{X}^{s}\pi^{s}\right)P^{T}=PF(\mathbf{Y},s)P^{T}$. Therefore, $\|P\mathbf{Y}P^{T}-\sum_{s}\lambda_{s}F(P\mathbf{Y}P^{T},s)\|_{\mathrm{Frob}}=\|\mathbf{Y}-\sum_{s}\lambda_{s}F(\mathbf{Y},s)\|_{\mathrm{Frob}}$. We refer to Section E.1 in the Suppl. Mat. for empirical demonstration.
@@
[…]
@@ 4. A Gradient Method for Gromov-Wasserstein Barycenters
In [15], given a network $\mathbb{X}$, the authors provide a way of computing directional derivatives of the functional $\mathbb{Y}\longmapsto GW^{2}(\mathbb{X},\mathbb{Y})$ in the direction of vectors in a ‘tangent space’ at $\mathbb{Y}$ [7, 14, 59]. Thus, in Subsections 4.1, 4.2, and 4.3 we shift to a different approach that allows to compute gradients of functionals in GW space using a blow-up technique, and thus defining a weaker version of GW barycenters as Karcher means. This culminates in our new contributions to the GW analysis problem (Section 4.4), including Theorem 4.7 and Algorithm 4. Connections between the fixed-point and gradient-based approaches are discussed in Section 4.5.
@@ 4.1. Tangent Space, Exponential Map, Directional Derivatives, and Gradients
As introduced in Section 2, let $\mathcal{GW}$ be the Gromov-Wasserstein space of equivalence classes of networks under weak isomorphism $\sim^{w}$. Thus, two networks $\mathbb{Y}$ and $\mathbb{Y}^{\prime}$ belong to the same class $\mathbb{Y}]$ if and only if $GW(\mathbb{Y},\mathbb{Y}^{\prime})=0$. In this case, a GW OT plan $\pi$ realizing $GW(\mathbb{Y},\mathbb{Y}^{\prime})=0$ is referred to as a self-coupling. In [15], the authors provide a formal tangent structure on the GW space $\mathcal{GW}$.In particular, they define the tangent space of $\mathcal{GW}$ at a network class $\mathbb{Y}]$ as the quotient space [^9] If we consider only the space of classes of mm-spaces or gm-spaces under the strong isomorphism defined in Definition 2.3, then each $L^{2}$-space in (24) is restricted to its subspace of symmetric functions.
(24) $\mathrm{Tan}_{[\mathbb{Y}]}:=\displaystyle\left(\bigcup_{(Y,\omega_{Y},\mu_{Y})\in[\mathbb{Y}]}L^{2}(Y\times Y,\mu_{Y}\otimes\mu_{Y})\right)/\simeq$
under the equivalence relation $\simeq$ defined as follows: for $(Y,\omega_{Y},\mu_{Y}),(Y^{\prime},\omega_{Y^{\prime}},\mu_{Y^{\prime}})\in[\mathbb{Y}]$, let $v\in L^{2}(Y\times Y,\mu_{Y}\otimes\mu_{Y})$, $v^{\prime}\in L^{2}(Y^{\prime}\times Y^{\prime},\mu_{Y^{\prime}}\otimes\mu_{Y^{\prime}})$, we say that $v\simeq v^{\prime}$ if and only if there is a self-coupling $\pi\in\Pi(\mu_{Y},\mu_{Y^{\prime}})$ such that
@@
We denote the class of $v$ as $[v]$. Each space $\mathrm{Tan}_{[\mathbb{Y}]}$ is equipped with the inner product
@@
(25) $\langle A,B\rangle_{L^{2}(Y\times Y,\mu_{Y}\otimes\mu_{Y})}=\sum_{i,j\in[M]}A_{ij}B_{ij}\mathrm{q}_{i}\mathrm{q}_{j}=:\mathrm{tr}_{\mathrm{q}}(A^{T}B),$
where $\mathrm{tr}_{\mathrm{q}}$ denotes the trace inner product weighted by $\mathrm{q}\otimes\mathrm{q}\in\mathbb{R}^{M\times M}$.To ensure that this defines a norm, we assume that $\mathrm{q}_{j}>0$ for all $j\in[M]$, meaning that the support of $\mu_{Y}$ consists precisely of the $M$ distinct points $\{y_{j}\}_{j=1}^{M}$, as required in Definition 2.1.
@@
(26) $v]\longmapsto[(Y,\omega_{Y}+v,\mu_{Y})]\qquad\text{ for }v\in L^{2}(Y\times Y,\mu_{Y}\otimes\mu_{Y}).$
Using our shorthand notation $(\mathbf{Y},\mathrm{q})\in\mathbb{R}^{M\times M}\times\mathcal{P}_{M}$ to denote a representative of $\mathbb{Y}]$, (26) can be identified with the map $v\mapsto(\mathbf{Y}+v,\mathrm{q})$ for all $v\in\mathbb{R}^{M\times M}$.
@@
Given a functional over the GW space, $G:\mathcal{GW}\to\mathbb{R}$, its directional derivative in the direction of $v]\in\mathrm{Tan}_{[\mathbb{Y}]}$ is defined by
(27) $D_{v}G([\mathbb{Y}]):=\lim_{t\searrow 0}\frac{G(Exp_{[\mathbb{Y}]}([tv]))-G([\mathbb{Y}])}{t},$
and its gradient [^10] If $G:\mathcal{GW}\to\mathbb{R}$ is differentiable, i.e, if all the directional derivatives (27) exist, and if, , its gradient at $\mathbb{Y}]$, as defined via Riesz representation (28), exists, then it is unique (see [15, Lem. 7] and the references therein). by a vector $\nabla G([\mathbb{Y}])\in\mathrm{Tan}_{[\mathbb{Y}]}$ satisfying
(28) $D_{v}G([\mathbb{Y}])=\langle[v],\nabla G([\mathbb{Y}])\rangle_{\mathrm{Tan}_{[\mathbb{Y}]}},\qquad\forall[v]\in\mathrm{Tan}_{[\mathbb{Y}]}.$
@@ 4.2. Blow-Ups
In [15], particular representatives of the equivalence classes under weak isomorphisms were considered using a blow-up technique, which we will describe. This blow-up idea avoids the need to fix a size $M$ in advance, as required in the approach of [49] and reviewed in this paper when studying the GW synthesis problem (7) in Section 3.1.
Figure 4. Sketch to visualize the classical barycentric projection. Gray dashed lines indicate the optimal transport plan $\pi$ between the source measure $\mu$ (black) and the target measure $\nu$ (blue). In general, if $\mu=\sum_{i}^{n}a_{i}\delta_{x_{i}}$ and $\nu=\sum_{j}^{m}b_{j}\delta_{y_{j}}$, then the barycentric projection measure (green) is supported at $\widehat{y_{i}}:=\frac{1}{a_{i}}\sum_{j}\pi_{ij}y_{j}$ and the mass at $\widehat{y_{i}}$ coincides with that of $x_{i}$ for all $i\in[n]$, thus, the optimal transport between the source measure and its barycentric projection is induced by a map (green dashed lines).
In a broader sense, this idea is closely related to the notion of barycentric projection widely used in linear OT [70] to recover transport maps from OT plans. While the barycentric projection effectively “collapses” mass (see Fig. 4), we will see that the blow-up technique instead “multiplies” the number of nodes (cf. Fig. 5 below). The parallelism is that, in both cases, one is led to consider optimal couplings that are induced by maps after obtaining distributions with the same number of particles and mass. For the GW case, we refer to the relevant work [6]. There, a single reference mm-space is fixed, and each object is embedded into a common linear space via a GW optimal coupling followed by a barycentric projection. This construction yields a local linear model of GW space and provides a computationally efficient approximation of the inherently nonlinear GW geometry. The focus of [6] is on defining a linear embedding of GW space and computing a single GW optimal coupling to a fixed reference, followed by a barycentric projection. Instead of embedding, we rely on the notion of weakly isomorphic networks: we consider multiple references and utilize the blow-up technique (which requires multiple optimal couplings) to obtain new weakly isomorphic networks with which we can compute “sums” and, moreover, “weighted averages”.
@@
Consider $(X=\{x_{i}\}_{i=1}^{N},\omega_{X},\mu_{X})$, $(Y=\{y_{j}\}_{j=1}^{M},\omega_{Y},\mu_{Y})$ finite networks, which we simply denote as $(\mathbf{X},\mathrm{p})\in\mathbb{R}^{N\times N}\times\mathcal{P}_{N}$, $(\mathbf{Y},\mathrm{q})\in\mathbb{R}^{M\times M}\times\mathcal{P}_{M}$ (we recall the notation $\omega_{X}(x_{i},x_{i^{\prime}})=\mathbf{X}[i,i^{\prime}]$, $\omega_{Y}(y_{j},y_{j^{\prime}})=\mathbf{Y}[j,j^{\prime}]$, $\mu_{X}=\sum_{i=1}^{N}\mathrm{p}[i]\delta_{x_{i}}$, $\mu_{Y}=\sum_{i=1}^{M}\mathrm{q}[j]\delta_{y_{j}}$). Let $\pi\in\Pi(\mathrm{p},\mathrm{q})$ be a coupling between $\mathrm{p}$ and $\mathrm{q}$. Consider the vectors $u({X})=(u(x_{1}),\dots u(x_{N}))$ and $v({Y})=(v(y_{1}),\dots,v(y_{M}))$ given by
@@
    Alignment: One can construct a (possibly enlarged) optimal coupling $\pi_{b}\in\Pi(\mathrm{p}_{b},\mathrm{q}_{b})$ for $GW(\mathbb{X}_{b},\mathbb{Y}_{b})$ that induces an optimal GW-transport map $T:X_{b}\to Y_{b}$, which is a bijection, i.e., a permutation of the indices in $M_{b}]$ (see [21] for existence of optimal GW maps in the context of mm-spaces under strong isomorphisms). As highlighted in [15], the crux of the blow-up construction relies on the observation that while $\mathbb{X}_{b}$ and $\mathbb{Y}_{b}$ maintain weak isomorphisms with $\mathbb{X}$ and $\mathbb{Y}$, respectively, the initial (optimal) transport plan $\pi$ between the original measures $\mu_{X},\mu_{Y}$ supported on the original nodes $X,Y$ (represented the vectors $\mathrm{p},\mathrm{q}$), lifts to a map T{T} between the new sets of nodes $X_{b},Y_{b}$ such that $T_{\#}\mu_{X_{b}}=\mu_{Y_{b}}$:
[…]
    For simplicity, after permuting and relabeling the nodes, if necessary, we will assume that $T$ is the identity map (i.e., we identify $X_{b}=Y_{b}$). Consequently, $\mathrm{p}_{b}=\mathrm{q}_{b}$, and $\pi_{b}$ can be regarded as the diagonal matrix with $\mathrm{q}_{b}$ as its diagonal entries, denoted as $\pi_{b}=\mathrm{diag}(\mathrm{q}_{b})$. See Fig. 5 for an illustration. Moreover,
    (29) $GW^{2}(\mathbb{X},\mathbb{Y})=\sum_{j,k\in[M_{b}]}|\mathbf{X}_{b}[j,k]-\mathbf{Y}_{b}[j,k]|^{2}\,\mathrm{q}_{b}[j]\mathrm{q}_{b}[k]=\mathrm{tr}_{\mathrm{q}_{b}}\left((\mathbf{X}_{b}-\mathbf{Y}_{b})^{T}(\mathbf{X}_{b}-\mathbf{Y}_{b})\right).$
@@
If we consider multiple templates $\{[\mathbb{X}^{s}]\}_{s=1}^{S}$ for $S\geq 1$, given a class $\mathbb{Y}]$, Algorithm 3 is a general blow-up, which facilitates the alignment of each template representative $\mathbb{X}^{s}=(X^{s},\omega_{X^{s}},\mu_{X^{s}})$ with a representative $\mathbb{Y}=(Y,\omega_{Y},\mu_{Y})$, ensuring that the new spaces $Y_{b}={X^{1}_{b}}=\dots={X^{S}_{b}}$ all contain the same number of nodes ($M_{b}$), share the same mass distribution $\mathrm{q}_{b}$ across the nodes, and the identity map is optimal for each GW problem between $\mathbb{X}^{s}_{b}$ and $\mathbb{Y}_{b}$ ($s\in[S]$). Note that the optimal plans between $\mathbb{X}^{s}_{b},\mathbb{X}^{r}_{b}$, for $s\neq r$, are not necessarily the identity. The network $\mathbb{Y}$ plays a special role in this algorithm, and thus when we apply it to templates $\{[\mathbb{X}^{s}]\}_{s=1}^{S}$, we say that we perform an optimal blow-up with respect to $\mathbb{Y}$.
@@
6:  $\pi$ $\leftarrow$ optimal plan for $GW((\mathbf{X},\mathrm{p}),(\mathbf{Y},\mathrm{q}))$
@@
Proposition 4.3 (Prop. 8 [15] - Revisited proof in Suppl. Mat. B.3).
@@
(30) $G_{\lambda}([\mathbb{Y}]):=\frac{1}{2}\sum_{s\in[S]}\lambda_{s}GW^{2}(\mathbb{X}^{s},\mathbb{Y}).$